Dozier Company produced and sold 1,000 units during its first month of operations. It reported the following costs and expenses for the month: Direct materials Direct labor Variable manufacturing overhead Fixed manufacturing overhead Total manufacturing overhead Variable Belling expense Fixed selling expense Total selling expense Variable administrative expense Fixed administrative expense Total administrative expense $ 89,000 $ 45,000 $23,000 34,000 $ 16,000 26,000 $ 6,000 $ 57,000 $ 42,000 $ 35,000 Required: 1. With respect to cost classifications for preparing financial statements: a. What is the total product cost? b. What is the total period cost? 2. With respect to cost classifications for assigning costs to cost objects: a. What is total direct manufacturing cost? b. What is the total indirect manufacturing cost? 3. With respect to cost classifications for manufacturers a. What is the total manufacturing cost? b. What is the total nonmanufacturing cost? c. What is the total conversion cost and prime cost? 4. With respect to cost classifications for predicting cost behavior a. What is the total variable manufacturing cost? b. What is the total fixed cost for the company as a whole? c. What is the variable cost per unit produced and sold? 5. With respect to cost classifications for decision.making: a. If Dozier had produced 1,001 units instead of 1,000 units, how much incremental manufacturing cost would it have incurred to make the additional unit?
Answer #1

Solution:

1 a. Total product cost = Direct material + Direct labor + Manufacturing OH costs

= $89,000 + 45,000 + 57,000 = $191,000

b. Total period cost = Selling Exp. + Administration Exp. = $42,000 + 35,000 = $77,000

2. a. Direct manufacturing cost = Direct material + Direct labor cost = $89,000 + 45,000

= $134,000

  b. Indirect manufacturing cost = Manufacturing OH = $23,000 + 34,000 = $57,000

3. a. Total manufacturing cost = $191,000

b. Total non manufacturing cost = $77,000

c. Total conversion = Direct labor + Manufacturing OH = $45,000 + 57,000 = $102,000

Prime cost = Direct material + direct labor =  $134,000

4. a.Total variable manufacturing cost = $89,000 + 45,000 +23,000= $157,000

b.Total Fixed costs for the company as a whole = Fixed manf. OH + Fixed selling exp. + Fixed Adm. Exp. = $34,000 + 26,000 + 29,000

= $89,000

c. Variable cost per unit produced = $157,000 (i.e 4a.) / 1,000 units = $157

  Variable cost per unit sold = ($157,000 + 16,000 +6,000) / 1,000 = $179

5 a. Incremental manufacturing cost = Excess units of production * Var. mfg. cost per unit

= 1 * $157 = $157